# Analysis of neutrino oscillation parameters in the light on quantum entanglement

### v1

@@ 3 Experimental setup and simulation details
Deep Underground Neutrino Experiment (DUNE) is one of the most promising upcoming long-baseline neutrino experiment. It has a baseline of 1300 km from the neutrino source at Fermi National Accelerator Laboratory (FNAL) to Sanford Underground Research Facility (SURF). For simulating the DUNE experiment, we use the files provided in the Technical Design Report paper DUNE:2021cuw . The files correspond to $624$ kt-MW-years of exposure: 6.5 years each in neutrino and antineutrino modes, with a 40 kt fiducial mass liquid argon time-projection chamber (LArTPC) far detector and a 120-GeV, 1.2 MW beam. This setup is equivalent to ten years of data collection based on the nominal staging assumptions described in DUNE:2020jqi .
@@
$33.41^{\circ}$ $8.58^{\circ}$ $42.2^{\circ}$ $7.41\times 10^{-5}$$\pm 2.507\times 10^{-3}$
Table 1: Oscillation parameters used to plot the probability and EOF, concurrence measurement.
@@ 4.1.1 Entanglement of formation
In the experimental setup of DUNE, as described in the previous section, we first verify entanglement in the framework of EOF, the theoretical aspects of which has been discussed in Section 2. The accelerator-based neutrino experiment predominantly generates muon-type neutrinos, establishing the initial flavor state as $\nu_{\mu}$ for this study. The goal is to quantify the entanglement between $\nu_{\mu}$ and the other two flavor states, $\nu_{e}$ and $\nu_{\tau}$. For the initial flavor state $\nu_{\mu}$, the EOF expression simplifies to,
@@
$\begin{split}&S_{\mu e}=-\frac{1}{2}\Big{[}P_{\mu e}\log_{2}P_{\mu e}+(P_{\mu\mu}+P_{\mu\tau})\log_{2}(P_{\mu\mu}+P_{\mu\tau})\Big{]},\\ &S_{\mu\tau}=-\frac{1}{2}\Big{[}P_{\mu\tau}\log_{2}P_{\mu\tau}+(P_{\mu\mu}+P_{\mu e})\log_{2}(P_{\mu\mu}+P_{\mu e})\Big{]},\\ &S_{\mu(e\tau)}=-\frac{1}{2}\Big{[}P_{\mu\mu}\log_{2}P_{\mu\mu}+(P_{\mu e}+P_{\mu\tau})\log_{2}(P_{\mu e}+P_{\mu\tau})\Big{]}.\\ \end{split}$ (16)
@@
From the panels, we can observe three main energy windows where the EOF is achieving two minima and one maximum. However, one of those two minima tends to zero and we define this minimum as global minimum. For the other minimum, we designate it as local minimum. The EOF global minimum is seen in the energy range ($1.0$ $1.5$) GeV, the maximum from $1.7$ GeV to $2.0$ GeV and the local minimum from $2.0$ GeV to $3.0$ GeV energy ranges. For our convenience, we describe these three energy windows as:
  • Energy window (EW1): $1.0-1.5$ GeV
  • Energy window (EW2): $1.7-2.0$ GeV
  • Energy window (EW3): $2.0-3.0$ GeV
@@
EW2: The energy window from $1.7$ GeV to $2.0$ GeV shows a peak in the total EOF value. The entanglement measure reaches its maximum when the appearance and disappearance probabilities of the bipartite system are equal, causing the system to behave as a maximally mixed state. Within this range, around $1.74$ GeV, we observe that both $P_{\mu\tau}$ and $P_{\mu\mu}$ equal to 0.5. At this point, $S_{\mu\tau}$ and $S_{\mu(e\tau)}$ reach their maximum values, indicating the emergence of maximal entanglement, particularly between $\nu_{\mu}$ and $\nu_{\tau}$. Theoretically, the EOF should peak at 1, representing maximal entanglement. However, due to the influence of $S_{\mu e}$, which contributes a background effect, the EOF slightly exceeds from 1. This suggests that while the system approaches maximal entanglement, the presence of $S_{\mu e}$ introduces additional contributions, deviating EOF from the theoretical maximum. In this energy range, all three flavor states are in maximally entangled state with a tendency of maximal mixing between $\nu_{\mu}$ and $\nu_{\tau}$, making it difficult to extract “pure” information about the oscillation parameters.
@@
EW3: At an energy of 2.5 GeV, we observe minima in both $S_{\mu\tau}$ and $S_{\mu(e\tau)}$. This can be physically interpreted as the entanglement measure reaching its lowest point when the appearance probability is at its maximum. This suggests that either the two-flavor states are nearly separable or the entanglement is not at its maximum. Within the $2.0-3.0$ GeV energy window, these minima indicate that the states are neither fully separable nor maximally entangled. The non-zero minima imply that, while the entanglement is not maximal, there is still a degree of entanglement present. The non-zero value confirms that some level of entanglement persists within the system. Additionally, at approximately $E_{\nu}=2.5$ GeV, the total EOF of the system also reaches a minimum. We interpret this minimum as a “local minimum” due to the non-zero entanglement value, which indicates that the system retains a degree of entanglement even at this energy.
It is noteworthy that $S_{\mu e}$ reaches a maximum at 2.5 GeV, which is counterintuitive given the expected trend for an entangled flavor state. This observation suggests a disentangling behavior of the $\nu_{e}$ state and supports the idea that $\nu_{e}$ can be considered a separable flavor state within the DUNE experimental framework. The entire analysis is conducted for $\delta_{CP}=0^{\circ}$, but similar behavior is observed for the maximum CP-violating phase ($\delta_{CP}=270^{\circ}$). The right column of the Fig. 1 shows the event rate distribution for $\nu_{e}$ appearance and $\nu_{\mu}$ disappearance. The lower panel of Figure 1 displays corresponding results for antineutrinos, which exhibit similar trends to those seen for neutrinos. To deepen our understanding, we will now examine entanglement in terms of concurrence.
@@ 4.1.2 Concurrence
By plotting these three terms individually, we can visualize the contribution of each flavor state to the total concurrence. This visualization aids in understanding the contributions from each flavor state and helps justify the monogamy inequality, which indicates strong correlations between flavor states and allows us to identify and characterize specific entangled neutrino flavors. In Figure 2, similar to Figure 1, the left column displays the concurrences as a function of energy along with the corresponding probabilities. Seven curves are shown: the red, green and purple curves represent the three concurrence terms, $C_{\mu e;\mu}^{2}$, $C_{\mu\tau;\mu}$ and $C^{2}_{\mu e\tau}$ respectively from Eq. 17, the blue curve shows the total concurrence, $C^{\mu^{2}}$ for the DUNE experimental setup, and the pink (orange) and cyan dashed curves represent the appearance (disappearance) probabilities. The analysis will focus on the three distinct energy windows discussed previously.
EW1: In the energy window of 1.01.51.01.51.0–1.51.0 – 1.5 GeV, we observe that the appearance probability $P_{\mu\mu}$ is 1, while $P_{\mu e}$ and $P_{\mu\tau}$ approaching to zero. This behavior indicates that the $\nu_{\mu}$ flavor state is completely disentangled from both $\nu_{e}$ and $\nu_{\tau}$, resulting in the three neutrino flavor states behaving as a completely separable system. This observation underscores the physical relevance of concurrence in our analysis, as it effectively captures the disentangled nature of the system within this energy range. Effectively it is noticeable in this energy window all terms in Eq. 17 approaches to zero which leads to disentangled state.
EW2: In the energy window between 1.7 and 2.0 GeV, there is a specific point, around 1.74 GeV, where $P_{\mu\tau}$ and $P_{\mu\mu}$ become equal. At this point, the total squared concurrence associated with tripartite neutrino system, $C^{\mu^{2}}$, reaches its maximum value, indicating a high degree of entanglement within the system. This maximum entanglement corresponds to a maximally mixed state, as evidenced by $P_{\mu\mu}=P_{\mu\tau}=0.5$. The plot shows that this maximal mixing is primarily due to the $\nu_{\mu}$ and $\nu_{\tau}$ states. Examining the individual contributions, we find that the entanglement measures $C^{2}_{\mu\tau;\mu}$ and $C^{2}_{\mu e;\mu}$ both reach peak values at a specific point in the energy spectrum. Notably, $C^{2}_{\mu\tau;\mu}$ exhibits a significantly higher amplitude compared to $C^{2}_{\mu e;\mu}$, indicating a stronger entanglement between the $\nu_{\mu}$ and $\nu_{\tau}$ states.
EW3: In the energy window 2.0–3.0 GeV, where we notice a dip in the values of $C^{2}_{\mu\tau;\mu}$, $C^{2}_{\mu e;\mu}$, and $C^{\mu^{2}}$. This dip suggests that the entanglement is reduced in this range, indicating a shift away from maximal entanglement. At approximately 2.5 GeV, the probability $P_{\mu\tau}$ reaches its peak, accompanied by a slight but noticeable increase in $P_{\mu e}$ and an almost vanishing value of $P_{\mu\mu}$. Despite this, $P_{\mu\tau}\neq 1$ at this energy, the flavor states are not fully disentangled; however, they do lose their maximally entangled state. This observation underscores that the nature of the entanglement is predominantly governed by the interaction between the dominant flavor states, $\nu_{\mu}$ and $\nu_{\tau}$. Based on our discussion of the concurrence, we conclude that in this energy window, the three-flavor neutrino system behaves as a bipartite state, with $\nu_{e}$ being in a separable state.
@@
Throughout the entire energy range, the term represented by the purple line, $4P_{\mu e}P_{\mu\tau}$, does not directly contribute to the physical interpretation of concurrence but influences the overall concurrence value. This term can be expressed as $\frac{C^{2}_{\mu e;\mu}C^{2}_{\mu\tau;\mu}}{4P^{2}_{\mu\mu}}$, representing the product of the squared concurrences of the $\nu_{\mu}-\nu_{e}$ and $\nu_{\mu}-\nu_{\tau}$ pairs. Despite its lack of direct physical significance in terms of concurrence, this term is crucial for validating the monogamy inequality of entanglement, which imposes constraints on how entanglement is distributed among multiple flavor states.
@@
In summary, by starting with the initial flavor state $\nu_{\mu}$, our analysis focuses on the bipartite entanglement between $\nu_{e}$ and $\nu_{\tau}$. Our findings demonstrate that, with $\nu_{\mu}$ as the initial flavor state, there is significant entanglement with $\nu_{\tau}$, including instances of maximal entanglement, while $\nu_{e}$ remains a separable state in certain energy ranges. This highlights the critical role of $\nu_{\mu}$ in the entanglement dynamics of the three-flavor system, particularly in its interactions with $\nu_{\tau}$.
@@ 4.2 Verification of Entanglement Monogamy
$\vspace{0cm}\begin{split}C^{2}_{\mu e;\mu}+C^{2}_{\mu\tau;\mu}\leq C^{\mu^{2}}\end{split}$ (18)
@@
In addition to the W state, there exists another significant entangled state in tripartite systems, namely the Greenberger-Horne-Zeilinger (GHZ) state, commonly represented as |$\ket{000}+\ket{111}$. The key difference between these two states lies in the behavior of entanglement under measurement. For the GHZ state, a measurement of any qubit in the |$\ket{111}$ state results in the destruction of entanglement, reflecting the nature of three-way entanglement D_r_2000 . In contrast, for the W state, a measurement on the first qubit still preserves the entanglement between the remaining two qubits, as seen in Eq. 19, indicating the presence of pairwise entanglement PhysRevA.101.032344 . This analysis suggests that, in the context of three-flavor neutrino oscillations, it is more likely to observe bipartite separable states rather than genuine tripartite entanglement. This experimental observation highlights the dominance of bipartite entanglement in neutrino flavor systems.
The W state is one of the two distinct classes of non-separable three-qubit states, the other being the GHZ state. These states cannot be transformed into each other via local quantum operations and are both considered maximally entangled for a three-party system. However, the W state exhibits a unique characteristic: if one of the three qubits is lost, the remaining two qubits retain their entanglement Swain:2021btc . This robustness of W-type entanglement is in stark contrast to the GHZ state, which becomes fully separable upon the loss of a single qubit Sk:2021dbf . The resilience of the W state under particle loss gives rise to the concept of distributed entanglement, which is quantified through a monogamy inequality known as the CKW inequality. This inequality was specifically formulated for the W class of states to describe the limits of entanglement sharing between qubits in a multipartite system. Both the W and GHZ states are fundamental to the advancement of quantum computing, providing essential resources for tasks such as quantum error correction and quantum communication protocols.

### v2

@@ 3 Experimental setup and simulation details
Deep Underground Neutrino Experiment (DUNE) is one of the most promising upcoming long-baseline neutrino experiment. It has a baseline of 1300 km from the neutrino source at Fermi National Accelerator Laboratory (FNAL) to Sanford Underground Research Facility (SURF). For simulating the DUNE experiment, we use the files provided in the Technical Design Report paper DUNE:2021cuw. The files correspond to $624$ kt-MW-years of exposure; 6.5 years each in neutrino and antineutrino modes, with a 40 kt fiducial mass liquid argon time-projection chamber (LArTPC) far detector and a 120-GeV, 1.2 MW beam. This setup is equivalent to ten years of data collection based on the nominal staging assumptions described in DUNE:2020jqi.
@@
$33.68^{\circ}$ $8.56^{\circ}$ $43.3^{\circ}$ $7.49\times 10^{-5}$ $\pm 2.513\times 10^{-3}$
Table 1: Oscillation parameters used to plot the probability and EOF, concurrence measurement Esteban:2024eli
@@ 4.1.1 Entanglement of formation
In the first hand we have used the entanglement of formation (EOF) as entanglement measures because from the mathematical expression of EOF we can see that it is directly related to the linear entropy of the individual components of the system which is the fundamental measure of a mixed system PhysRevA.67.022110. Hence using EOF as a entanglement measures justifiably signifies the mixedness within the neutrino system and the amount of entanglement present in it. In the experimental setup of DUNE, as described in the previous section, we first verify entanglement in the framework of EOF, the theoretical aspects of which has been discussed in Section 2. The accelerator-based neutrino experiment predominantly generates muon-type neutrinos, establishing the initial flavor state as $\nu_{\mu}$ for this study. The goal is to quantify the entanglement between $\nu_{\mu}$ and the other two flavor states, $\nu_{e}$ and $\nu_{\tau}$. For the initial flavor state $\nu_{\mu}$, the EOF expression simplifies to,
@@
[…]
@@
From the panels, we can observe three main energy windows where the EOF is achieving two minima and one maximum. However, one of those two minima tends to zero and we define this minimum as global minimum. For the other minimum, we designate it as local minimum. The EOF global minimum is seen in the energy range ($1.0$ - $1.5$) GeV, the maximum from $1.5$ GeV to $2.0$ GeV and the local minimum from $2.0$ GeV to $3.0$ GeV energy ranges. For our convenience, we describe these three energy windows as:
  • Energy window (EW1): 1.01.51.0-1.5 GeV
  • Energy window (EW2): 1.52.01.5-2.0 GeV
  • Energy window (EW3): 2.03.02.0-3.0 GeV
@@
EW2: The energy window from $1.5$ GeV to $2.0$ GeV shows a peak in the total EOF value. The entanglement measure reaches its maximum when the appearance and disappearance probabilities of the bipartite system are equal, causing the system to behave as a maximally mixed state. Within this range, around $1.74$ GeV, we observe that both $P_{\mu\tau}$ and $P_{\mu\mu}$ equal to 0.5. At this point, $S_{\mu\tau}$ and $S_{\mu(e\tau)}$ reach their maximum values, indicating the emergence of maximal entanglement, particularly between $\nu_{\mu}$ and $\nu_{\tau}$. The upper bound of the EOF is given by $\log_{2}d$, where $d$ denotes the dimension of the Hilbert space of each subsystem Bennett:1996gf. For a two-qubit system $(d=2)$, the maximum EOF is $\log_{2}2=1$. In contrast, for neutrinos, where $d=3$, the upper limit becomes $\log_{2}3\approx 1.58$, which exceeds 1. Hence, obtaining EOF values greater than 1 in this case is physically consistent. In this energy range, all three flavor states are in maximally entangled state with a tendency of maximal mixing between $\nu_{\mu}$ and $\nu_{\tau}$, making it difficult to extract “pure” information about the oscillation parameters.
@@
EW3: At an energy of 2.5 GeV, we observe minima in both $S_{\mu\tau}$ and $S_{\mu(e\tau)}$. This can be physically interpreted as the entanglement measure reaching its lowest point when the appearance probability is at its maximum. This suggests that either the two-flavor states are nearly separable or the entanglement is not at its maximum. Within the 2.03.02.0-3.0 GeV energy window, these minima indicate that the states are neither fully separable nor maximally entangled. The non-zero minima imply that, while the entanglement is not maximal, there is still a degree of entanglement present. The non-zero value confirms that some level of entanglement persists within the system. Additionally, at approximately $E_{\nu}=2.5$ GeV, the total EOF of the system also reaches a minimum. We interpret this minimum as a “local minimum” due to the non-zero entanglement value, which indicates that the system retains a degree of entanglement even at this energy. The lower panel of Figure 1 displays corresponding results for antineutrinos, which exhibit similar trends to those seen for neutrinos.
Although, the EOF provides an interesting aspects in the entanglement measures in a mixed quantum system, it is very hard to compute. As EOF is a logarithmic function, it is very difficult to analyze the nature of the entanglement (bipartite or pure tripartite entanglement) within the system. It is noteworthy that $S_{\mu e}$ reaches a maximum at 2.5 GeV, which is counterintuitive given the expected trend for an entangled flavor state which is hard to explain.
The above analysis suggests clearly that although it depicts a hint of maximal entanglement between muon and tau flavor state of neutrino but due to logarithmic nature of EOF, we can not conclusively say which flavor states are entangled. Hence we proceed further to analyse with respect to examine the nature of entanglement in terms of squared-concurrence.
@@ 4.1.2 Concurrence
By plotting these three terms individually, we can visualize the contribution of each flavor state to the total concurrence. This visualization aids in understanding the contributions from each flavor state and helps to justify the monogamy inequality, which indicates strong correlations between flavor states and allows us to identify and characterize specific entangled neutrino flavors.
In Fig. 2, the left column displays the concurrences as a function of energy along with the corresponding probabilities. In order to generate the plots, we have taken the values of neutrino oscillation parameters from Table 1 with $\delta_{CP}$ as $0^{\circ}$. Seven curves are shown: the red, purple and green curves represent the three concurrence terms, $C_{\mu e;\mu}^{2}$, $C^{2}_{\mu\tau;\mu}$ and $C^{2}_{\mu e\tau}$ respectively from Eq. 17, the blue curve shows the total concurrence, $C^{\mu^{2}}$ for the DUNE experimental setup, and the pink (orange) and cyan dashed curves represent the appearance (disappearance) probabilities. The analysis will focus on the three distinct energy windows discussed previously.
EW1: In the energy window of 1.01.51.0–1.5 GeV, we observe that the appearance probability $P_{\mu\mu}$ is 1, while $P_{\mu e}$ and $P_{\mu\tau}$ approaching to zero. This behavior indicates that the $\nu_{\mu}$ flavor state is completely disentangled from both $\nu_{e}$ and $\nu_{\tau}$, resulting in the three neutrino flavor states behaving as a completely separable system. This observation underscores the physical relevance of concurrence in our analysis, as it effectively captures the disentangled nature of the system within this energy range. Effectively it is noticeable in this energy window all terms in Eq. 17 approaches to zero which leads to disentangled state.
EW2: In the energy window between 1.5 and 2.0 GeV, there is a specific point, around 1.74 GeV, where $P_{\mu\tau}$ and $P_{\mu\mu}$ become equal. At this point, the total squared concurrence associated with tripartite neutrino system, $C^{\mu^{2}}$, reaches its maximum value, indicating a high degree of entanglement within the system. This maximum entanglement corresponds to a maximally mixed state, as evidenced by $P_{\mu\mu}=P_{\mu\tau}=0.5$. The plot shows that this maximal mixing is primarily due to the $\nu_{\mu}$ and $\nu_{\tau}$ states. Examining the individual contributions, we find that the entanglement measures $C^{2}_{\mu\tau;\mu}$ and $C^{2}_{\mu e;\mu}$ both reach peak values at a specific point in the energy spectrum. Notably, $C^{2}_{\mu\tau;\mu}$ exhibits a significantly higher amplitude compared to $C^{2}_{\mu e;\mu}$, indicating a stronger entanglement between the $\nu_{\mu}$ and $\nu_{\tau}$ states. It is interesting to note that, the total concurrence value exceeds the value 1. In our analysis, $C^{\mu^{2}}$ represents the total concurrence, not the concurrence for just a single bi-partition. For each bi-partition, the concurrence is still less than or equal to one. When we add the three concurrences for three different bi-partitions, it can exceed value 1. The upper bound on total $C^{\mu^{2}}$ should be $\frac{2(d-1)}{d}$ with $d$ being the dimension of the Hilbert space of each subsystem Li:2005mbb; Dur:2000zz. For instance, in the case of two qubits (d= 2), the maximum possible value of $C^{\mu^{2}}$ is 1. However, for three-dimensional systems such as neutrinos (d=3), the upper bound increases to $\frac{2\times(3-1)}{3}=\frac{4}{3}$ which is greater than 1. Consequently, the total concurrence in our results correctly attains this maximum value of $\frac{4}{3}=1.33$.
EW3: In the energy window 2.0–3.0 GeV, where we notice a dip in the values of $C^{2}_{\mu\tau;\mu}$, $C^{2}_{\mu e;\mu}$, and $C^{\mu^{2}}$. This dip suggests that the entanglement is reduced in this range, indicating a shift away from maximal entanglement.
At approximately 2.5 GeV, the probability $P_{\mu\tau}$ reaches its peak, accompanied by a slight but noticeable increase in $P_{\mu e}$ and an almost vanishing value of $P_{\mu\mu}$. Despite this, $P_{\mu\tau}\neq 1$ at this energy, the flavor states are not fully disentangled; however, they do lose their maximally entangled state. This observation underscores that the nature of the entanglement is predominantly governed by the interaction between the dominant flavor states, $\nu_{\mu}$ and $\nu_{\tau}$. Based on our discussion of the concurrence, we conclude that in this energy window, the three-flavor neutrino system behaves as a bipartite state, with $\nu_{e}$ being in a separable state.
@@
Throughout the entire energy range, the term represented by the green line, $4P_{\mu e}P_{\mu\tau}$, does not directly contribute to the physical interpretation of concurrence but influences the overall concurrence value. This term can be expressed as $\frac{C^{2}_{\mu e;\mu}C^{2}_{\mu\tau;\mu}}{4P^{2}_{\mu\mu}}$, representing the product of the squared concurrences of the $\nu_{\mu}-\nu_{e}$ and $\nu_{\mu}-\nu_{\tau}$ pairs. Despite its lack of direct physical significance in terms of concurrence, this term is crucial for validating the monogamy inequality of entanglement, which imposes constraints on how entanglement is distributed among multiple flavor states.
Figure 3: Left (right) panel shows the oscillogram plot of EOF (concurrence) at three different energy windows for whole $\delta_{CP}$ range.
@@
In summary, by starting with the initial flavor state $\nu_{\mu}$, our analysis focuses on the bipartite entanglement between $\nu_{e}$ and $\nu_{\tau}$. Our findings demonstrate that, with $\nu_{\mu}$ as the initial flavor state, there is significant entanglement with $\nu_{\tau}$, including instances of maximal entanglement, while $\nu_{e}$ remains a separable state in certain energy ranges. This highlights the critical role of $\nu_{\mu}$ in the entanglement dynamics of the three-flavor system, particularly in its correlation with $\nu_{\tau}$.
For the analysis of EOF and concurrence, we have used a fixed value of $\delta_{CP}=0^{\circ}$. However, the general behavior of the entanglement measure plots as a function of neutrino energy remains consistent for other values of $\delta_{CP}$. To examine the dependency of EOF and concurrence on $\delta_{CP}$, an oscillogram is presented in Figure 3. The left panel of the figure illustrates the variation of EOF, while the right panel shows the variation of concurrence for the three previously discussed energy windows across different $\delta_{CP}$ values. In both panels, dark green represents the minimum value of the entanglement measure, while dark pink indicates the maximum values of EOF and concurrence. Both panels reveal that around 1.27 GeV, a global minimum occurs, at 1.74 GeV, a maximum is observed, and at 2.5 GeV, a local minimum appears. This behavior remains consistent across the entire range of $\delta_{CP}$ values. Therefore, although $\delta_{CP}=0^{\circ}$ was used in Figures 1 and 2, the result remains similar for other values of $\delta_{CP}$ as well.
@@ 4.2 Verification of Entanglement Monogamy
$\vskip 0.0pt\begin{split}C^{2}_{\mu e;\mu}+C^{2}_{\mu\tau;\mu}\leq C^{\mu^{2}}\end{split}$ (18)
@@
[…]